# Benford behavior resulting from stick and box fragmentation processes

### v2

@@
Benford behavior resulting from stick and box fragmentation processes
Bruce Fang fangbaojun2002@gmail.com Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267  and  Steven J. Miller sjm1@williams.edu, Steven.Miller.MC.96@aya.yale.edu Department of Mathematics and Statistics, Williams College, Williamstown, MA 01267
(Date: August 19, 2025)
@@
Benford’s law is the statement that in many real-world data set, the probability of having digit $d$ in base $B$, where $1\leq d\leq B$, as the first digit is $\log_{B}\left((d+1)/d\right)$. We sometimes refer to this as weak Benford behavior, and we say that a data set exhibits strong Benford behavior in base $B$ if the probability of having significand at most $s$, where $s\in[1,B)$, is $\log_{B}(s)$. We examine Benford behaviors in two different probabilistic model: stick and box fragmentation. Building on the joint work of [B–] on the 1-dimensional stick fragmentation model, we employ combinatorial identities on multinomial coefficients to reduce the high dimensional stick fragmentation model to the 1-dimensional model and provide a necessary and sufficient condition for the lengths of the stick fragments to converge to strong Benford behavior. Then we answer a conjecture posed by [BDMMM] on the high dimensional box fragmentation model. Using tools from order statistics, we prove that under some mild conditions, faces of any arbitrary dimension of the box have total volume converging to strong Benford’s behavior.
1. Introduction
@@
At the beginning of the $20$th century, astronomer and mathematician Simon Newcomb [Ne] noticed an unusual pattern in the logarithmic tables he used at work. The early pages of the books were far more worn than the later ones, suggesting that numbers starting with smaller digits were consulted more often. From this, Newcomb inferred that people had a tendency to encounter numbers beginning with $1$ more frequently. Specifically, he found that $1$ appeared as the leading digit about $30\%$ of the time, $2$ about $17\%$, with the frequency decreasing for larger digits. Although he formulated a mathematical explanation for this curious phenomenon, his discovery initially went largely overlooked.
@@
It took another 57 years after Newcomb’s discovery for physicist Frank Benford to make the exact same observation as Newcomb: the first pages of logarithmic tables were used far more than others. He formulated this law as follows.
Definition 1.1.
@@
[Ben, Page 554] Data exhibits (weak) Benford behavior base $B$ if the frequency $F_{d}$ of leading digit $d$ is
$F_{d}\ =\ \log_{B}\left(\frac{d+1}{d}\right).$ (1.1)
@@
Nowadays, Benford’s Law is used in detecting many different forms of fraud, and its prevalence in the world fascinates not only mathematicians, but many other scientists as well (to learn more about Benford’s law and its many applications, we recommend [BeHi, Nig, Mil1, NiMi] to name a few).
@@
In 1986, Lemons [Lem] proposed using Benford’s law to analyze the partitioning of a conserved quantity. Since then, driven by the potential application to nuclear fragmentation, mathematicians and physicists have taken an interest in the Benford behavior of various fragmentation processes. Among these processes of interest is stick fragmentation. In a 1-dimensional stick fragmentation model, one begins with a stick of length $L$. Draw $p_{1}$ from a probability distribution on $(0,1)$. This fragments the stick into two sub-sticks of lengths $p_{1}L$ and $(1-p_{1})L$. For each sub-stick, draw two probabilities $p_{2}$ and $p_{3}$ from the same distribution. We perform this fragmentation process for a total of $N$ stages, where at each stage we fragment all the sub-sticks created during the previous stage according to probabilities. Hence, by the end of $N$ stages, there are $2^{N}$ sticks. Of particular interest is whether or not this fragmentation process converges to Benford’s behavior.
1.1. Previous Work on Fragmentation
@@
An important definition when studying a more precise statistical version of Benford behavior is the notion of the significand of a real number, i.e., its leading digits in scientific notation.
Definition 1.2.
Given a positive real number $x$, we say that its significand base $B>1$, denoted $S_{B}(x)$, is the unique real number $S_{B}(x)\in[1,B)$ such that $x=S_{B}(x)\cdot B^{k}$, where $k$ is an integer.
As is common practice with these techniques involving proofs of Benford behavior, we define a more general version of Benford behavior that allows for processes that do not necessarily exhibit Benford behavior at first but converge to it in the limit.
Definition 1.3.
We say that a sequence of random variables $X^{(n)}$ converges to strong Benford behavior base $B$ if and only if
$\mathbb{P}(S_{B}(X^{(n)})\leq s)\to\log_{B}(s)$ (1.2)
for all $s\in[1,B]$. Notice by compactness that this implies uniform convergence of (1.2).
An equivalent formulation to the above is the Uniform Distribution Characterization, which is especially suited for investigation of products of random variables; see [Dia] for a proof.
Proposition 1.4 (Uniform Distribution Characterization).
@@
[Dia] A sequence of random variables $X^{(n)}$ converges to strong Benford behavior in base $B$ if and only if
$\mathbb{P}\left(\log_{B}(X^{(n)})\textup{ mod }1\leq t\right)\rightarrow t,$ (1.3)
for all $t\in[0,1]$. If (1.4) is satisfied, then we say that $\log_{B}(X^{(n)})$ converges to being equidistributed mod 1.
We may now state the previous results on stick fragmentation. Becker et al. [B–] proved a theorem regarding fixed proportion 1-dimensional stick fragmentation (compare with their Theorem 1.11).
Theorem 1.5.
@@
[B–] Choose any $p\in(0,1)$. In Stage 1, cut a given stick at proportion $p$ to create two pieces of length $p$ and 1p1-p. In Stage 2, cut each resulting piece into two pieces at the same proportion $p$. Perform this process for a total of $N$ stages, generating $2^{N}$ sticks with $N+1$ distinct lengths (assuming $p\neq 1/2$) given by
@@
where the frequency of $x_{n}$ is $\binom{N}{n}/2^{N}$. Choose $y$ so that $B^{y}=(1-p)/p$, which is the ratio of adjacent lengths (i.e., $x_{i+1}/x_{i}$). The decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y\not\in\mathbb{Q}$. If $y$ has finite irrationality exponent, the convergence rate can be quantified in terms of that exponent, and there is a power savings.
Theorem 1.5 suggests the possibility of examining another fixed proportion 1-dimensional stick fragmentation model, where we cut a piece into an arbitrary $m\geq 2$ number of pieces according m1m-1 fixed proportions at each stage. Using new techniques involving multinomial coefficients to reduce the problem into the original fixed proportion 1-dimensional stick fragmentation model in [B–], this paper establishes the following result (see Section 2).
Theorem 1.6.
@@
For any integer $m>2$, choose $p_{1},p_{2},\dots,p_{m-1}\in(0,1)$ such that $p_{1}+p_{2}+\dots+p_{m-1}<1$. Set $p_{m}:=1-(p_{1}+p_{2}+\dots+p_{m-1})$. At each stage, we cut all sticks according to proportions $p_{1},p_{2},\dots,p_{m-1}$ to create $m$ pieces. After stage $N$, we have $m^{N}$ sticks in total, of lengths
$A_{k_{1},k_{2},\dots,k_{m}}\ =\ p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}},$ (1.5)
for $0\leq k_{1},k_{2},\dots,k_{m}\leq N$ such that $k_{1}+k_{2}+\cdots+k_{m}=N$. Let $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for $1\leq i\leq m-1$. Then the decomposition process results in stick lengths that converge to strong Benford’s behavior base $B$ if and only if $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$.
Another possibility is to consider fragmentation processes in higher dimensions. This was considered by Durmić and Miller in [DM] and Benford behavior was established in the case of volume of an arbitrary $m$-dimensional box under mild assumptions. This was generalized by Betti et al. in [BDMMM] to any $d$-dimensional volume of $d$-dimensional faces of an arbitrary $m$-dimensional box, where $d\leq m$. To state their results, we start with some definitions.
Definition 1.7.
@@
We say that a set $\mathfrak{B}\subset\mathbb{R}^{m}$ is an $m$-dimensional box if it is a set of the form $a_{1},b_{1}]\times\cdots\times[a_{m},b_{m}]\subset\mathbb{R}^{m}$, where $a_{i}<b_{i}$ are finite numbers.
Definition 1.8.
@@
$\textup{Vol}_{d}(\mathfrak{B})\ :=\ 2^{m-d}\sum_{|I|=d}\prod_{i\in I}(b_{i}-a_{i}),$ (1.6)
@@
[BDMMM] established a sufficient condition for strong Benford behavior to emerge (see their Theorem 1.9), which involves the maximum $d$-dimensional volume of a $d$-dimensional face.
Theorem 1.10 (Maximum Criterion).
@@
[BDMMM] Let $\mathfrak{B}:=\mathfrak{B}_{0}$ be a fixed $m$-dimensional box and $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\cdots$ be a linear-fragmentation process whose proportion cuts $P_{i}$ have probability density functions $f_{i}:(0,1)\rightarrow(0,\infty)$. Let
$V_{d}^{(N)}\ :=\ \textup{Vol}_{d}(\mathfrak{B}_{N})$ (1.7)
@@
be the sequence of volumes obtained from this process and $\mathfrak{m}_{d}^{(N)}$ denote the maximum product of $d$ sides of $\mathfrak{B}_{N}$. If $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$, then so too does $V_{d}^{(N)}$ converge to strong Benford behavior base $B$ as $N\rightarrow\infty$.
In the same paper, the Maximum Criterion was verified for $d=m$ for any linear fragmentation process, and for $d=1$ in a special case.
Theorem 1.11.
@@
Let $P_{i}^{(j)}$ be i.i.d. log-uniform distributions, i.e., $\log_{B}P_{i}^{(j)}$’s have uniform distributions. In the case of $d=1$, i.e., perimeter, the maximum side length of each box
$\mathfrak{m}_{1}^{(N)}\ =\ \textup{max}_{1\leq i\leq m}P_{i}^{(1)}\cdots P_{i}^{(N)}$ (1.8)
converges to strong Benford behavior base $B$ as $N\rightarrow\infty$.
@@
Hence, it was conjectured in the same paper (see Conjecture 4.1) that for any linear fragmentation process, $m_{d}^{(N)}$ converges to strong Benford behavior.
Conjecture 1.12.
[BDMMM] Every linear fragmentation process satisfies the Maximum Criterion in all dimensions $1\leq d\leq m$.
@@
We prove the conjecture for any $m\geq 2$ and $d\leq m$ under some mild conditions on the linear fragmentation process (see Section 3).
Theorem 1.13.
@@
Suppose that $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\cdots$ is a linear fragmentation process on an $m$-dimensional box for an arbitrary $m\geq 2$, such that the proportion cut $P_{i}$ at every stage satisfies the following.
  1. (1)
    We have $\mathbb{E}[\log_{B}(P_{i})]=0$ and $\textup{Var}[\log_{B}(P_{i})]=1$.
  2. (2)
    We have $\log(P_{i})$ is supported on [C,C][-C,C], where $C>0$ is a constant.
  3. (3)
    Suppose that $P^{(k)}_{i}$ is a sequence of i.i.d. random variables $\sim P_{i}$. Define the random variable
    $Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (1.9)
    Then $Z_{i}^{(N)}$ has probability density function $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$ and cumulative density function $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\varphi(x)$ and $\Phi(x)$ are the probability density function and cumulative density function of the standard normal $N(0,1)$, $A(x)=O(N^{-1/2-\delta})$ and $B(x)=O(N^{-\delta})$ for some $\delta>0$.
@@
Then this linear fragmentation process satisfies the condition of the Maximum Criterion for any $d\leq m$, i.e., $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$, thus by Theorem 1.10, $V_{d}^{(N)}$ converges to strong Benford behavior base $B$ as $N\rightarrow\infty$ for any $d\leq m$.
2. Proof of Theorem 1.6
2.1. Preliminary

### v3

@@
Abstract
@@
Benford’s law is the statement that in many real-world data set, the probability of having digit $d$ in base $B$ as the first digit is $\log_{B}\left((d+1)/d\right)$ for all $1\leq d\leq B$. We sometimes refer to this as weak Benford behavior, and we say that a data set satisfies strong Benford behavior in base $B$ if the probability of having significand at most $s$ is $\log_{B}(s)$ for all $s\in[1,B)$. We examine Benford behaviors in two different probabilistic model: stick and box fragmentation models. Building on the joint work of Becker et al. [1] on the single proportion stick fragmentation model, we employ combinatorial identities on multinomial coefficients to reduce the multi-proportion stick fragmentation model to the single proportion model. We then provide a necessary and sufficient condition for the lengths of the stick fragments to converge to strong Benford behavior along with a quantification of the discrepancy from uniform distribution on $[0,1]$ in terms of irrationality exponent. Then we answer a conjecture posed by Betti et al. [6] on the high dimensional box fragmentation model. Using tools from Fourier analysis and order statistics, we prove that under some conditions, faces of any arbitrary dimension of the box have total volume converging to strong Benford’s behavior.
keywords:
Benford’s law , high-dimensional fragmentation , equidistribution mod 1 , irrationality exponent , multinomial identities , order statistics
2008 MSC:
60A10 , 11K06 (primary) , 60E10 (secondary)
journal: Journal of Number Theory
[…]
[inst1]organization=Department of Mathemtics, addressline=Williams College, city=Williamstown, postcode=01267, state=MA, country=USA
1 Introduction
@@
At the end of the 19th century, astronomer and mathematician Simon Newcomb [18] noticed an unusual pattern in the logarithmic tables he used at work. The early pages of the books were far more worn than the later ones, suggesting that numbers starting with smaller digits were consulted more often. Specifically, Newcomb found that $1$ appeared as the leading digit for about $30\%$ of the time, $2$ for about $17\%$, with the frequency decreasing for larger digits. Although he formulated a mathematical law for this curious phenomenon, his discovery initially went largely overlooked.
@@
It took over 50 years after Newcomb’s discovery for physicist Frank Benford to make the same observation as Newcomb. He formulated this law as follows. Although the law was originally only stated for base 10, we state the generalized form for any base $B\geq 2$.
Definition 1.
@@
[5, Page 554] A data set is said to satisfy Benford’s law for the leading digit in base $B$ if the frequency $F_{d}$ of leading digit $d$ is
$F_{d}\ =\ \log_{B}\left(\frac{d+1}{d}\right).$ (1)
@@
Nowadays, Benford’s Law is used in detecting many different forms of fraud, and its prevalence in the world fascinates not only mathematicians, but many other scientists as well (to learn more about Benford’s law and its many applications, we recommend [3, 17, 19] to name a few).
@@
In 1986, Lemons [16] proposed applying Benford’s law to the analysis of the partitioning of a conserved quantity. Motivated in part by potential applications to nuclear fragmentation, both mathematicians and physicists have since investigated the Benford behavior of a range of fragmentation models. One such model is stick fragmentation. In a 1-dimensional stick fragmentation model, we begin with a stick of length $L$. Draw $p_{1}$ from a probability distribution on $(0,1)$. This divides the stick into two sub-sticks of lengths $p_{1}L$ and $(1-p_{1})L$. For each resulting sub-stick, draw two probabilities $p_{2}$ and $p_{3}$ from the same distribution and fragment them accordingly. This procedure is repeated for a total of $N$ stages, with every sub-stick produced at a given stage being fragmented at the next stage using newly drawn probabilities. After $N$ stages, the process yields $2^{N}$ sticks. A central question is whether the distribution of stick lengths produced by this process converges to Benford’s behavior.
1.1 Previous Work on Fragmentation
@@
An key definition in formulating a more general version of Benford behavior is the notion of the significand of a real number, i.e., its leading digits in scientific notation.
Definition 2.
For any $x>0$, we can uniquely write
$x\ =\ S_{B}(x)\cdot B^{k_{B}(x)},$ (2)
where $S_{B}(x)\in[1,B)$ and $k_{B}(x)\in\mathbb{Z}$. Equivalently, $k_{B}(x)=\lfloor\log_{B}(x)\rfloor$ and $S_{B}(x)=x/B^{k_{B}(x)}$. We call $S_{B}(x)$ the significand of $x$.
We now define a more general version of Benford behavior that captures processes that only display Benford behaviors in the limit.
Definition 3.
@@
A sequence of random variables $\{X^{(N)}\}_{N=1}^{\infty}$ is said to converge to strong Benford behavior in base $B$ if
$\lim_{N\rightarrow\infty}\mathbb{P}(S_{B}(X^{(N)})\leq s)\ =\ \log_{B}(s)$ (3)
for all $s\in[1,B)$..
An equivalent formulation to the above is the Uniform Distribution Characterization, which is especially suited for investigation of products of random variables; see [8] for a proof.
Proposition 1 (Uniform Distribution Characterization).
[8] A sequence of random variables $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior in base $B$ if and only if
$\lim_{N\rightarrow\infty}\mathbb{P}\left(\log_{B}(X^{(N)})\textup{ mod }1\leq t\right)\ =\ t,$ (4)
for all $t\in[0,1)$. If (4) is satisfied, then we say that $\{\log_{B}(X^{(N)})\}_{N=1}^{\infty}$ converges to being equidistributed mod 1.
We are now ready to review some previous results on stick fragmentation models and introduce the two models of our paper. The first model is an extension of the model introduced by Becker et al. called the fixed single proportion stick fragmentation model [1]. Their main result on this model is a necessary and sufficient condition for the distribution of stick lengths to converge to strong Benford behavior as well as a quantification of the error term in terms of irrationality exponent, which we define now.
Definition 4.
[15] Suppose that $x$ is a real number. The irrationality exponent $\mu$ of $x$ is the supremum of the set of $\mu$ such that $0<|x-p/q|<1/q^{\mu}$ is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $q>0$. If such a set does not exist, then we say $x$ has irrationality exponent $\infty$.
Definition 5.
[15] For a finite sequence $\{x_{i}\}_{i=1}^{N}$, denote by $A([a,b),N)$ the number of $x_{i}$’s such that $x_{i}\textup{ mod 1}\in[a,b)$ for $0\leq a<b\leq 1$. Then we call the number
$D_{N}\ :=\ \sup_{0\leq a<b\leq 1}\left|\frac{A([a,b),N)}{N}-(b-a)\right|$ (5)
the discrepancy of the sequence. It measures how close a sequence is from being uniform mod 1.
Now we are ready to state a main theorem of [1].
Theorem 1.
@@
[1] Fix a proportion $p\in(0,1)$ and a stick of length $L$. At Stage 1, cut the stick at proportion $p$ to create two sub-sticks of lengths $Lp$ and L(1p)L(1-p). At Stage 2, cut each of these two sub-sticks into two smaller sub-sticks at the same proportion $p$. Repeat this procedure for a total of $N$ stages, generating $2^{N}$ sticks with $N+1$ distinct lengths (assuming $p\neq 1/2$) given by
@@
where the frequency of $x_{n}$ is $\binom{N}{n}/2^{N}$. Consider the sequence of random lengths $\{X^{(N)}\}_{N=1}^{\infty}$, where $X^{(N)}$ is the length of a stick picked uniformly at random from the $2^{N}$ sticks obtained at stage $N$. Choose $y$ so that $B^{y}=(1-p)/p$, which is the ratio of adjacent lengths (i.e., $x_{i+1}/x_{i}$). Then $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior base in $B$ if and only if $y\not\in\mathbb{Q}$. If $y$ has a finite irrationality exponent, then for sufficiently large $N$, the discrepancy of the sequence of the $2^{N}$ stick lengths (not distinct) can be quantified in terms of that exponent, and there is a power savings.
Theorem 1 suggests the possibility of generalizing the fixed single proportion stick fragmentation model to a fixed multi-proportion stick fragmentation model, where at each stage, instead of cutting a stick into 2 sub-sticks according to a fixed proportion, we cut a stick into an arbitrary but fixed $m$ number of sub-sticks according to m1m-1 fixed proportions. Again, we are interested in the conditions under which the distribution of stick lengths produced in this model converges to strong Benford’s behavior. A main result of this paper provides a definitive answer to this question, as follows:
Theorem 2.
@@
For a fixed integer $m>2$, choose $p_{1},p_{2},\dots,p_{m-1}\in(0,1)$ such that $p_{1}+p_{2}+\dots+p_{m-1}<1$. Set $p_{m}:=1-(p_{1}+p_{2}+\dots+p_{m-1})$. Now, consider a stick of length $L$. At each stage, we cut every remaining stick into $m$ sub-sticks according to proportions $p_{1},p_{2},\dots,p_{m-1}$. After stage $N$, we have $m^{N}$ sticks in total, of lengths
$A^{(N)}_{k_{1},k_{2},\dots,k_{m}}\ =\ Lp_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{m}^{k_{m}},$ (7)
for all $0\leq k_{1},k_{2},\dots,k_{m}\leq N$ such that $k_{1}+k_{2}+\cdots+k_{m}=N$. Consider the sequence of random lengths $\{X^{(N)}\}_{N=1}^{\infty}$, where $X^{(N)}$ is the length of a stick picked uniformly at random from the $m^{N}$ sticks obtained at stage $N$. Let $y_{i}=\log_{B}(p_{i}/p_{i+1})$ for all $1\leq i\leq m-1$. Then $\{X^{(N)}\}_{N=1}^{\infty}$ converges to strong Benford behavior base in $B$ if and only if $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$. Let $\kappa_{0}$ be the least irrationality exponent among all the irrational $y_{i}$’s. Then for sufficiently large $N$, the discrepancy of the sequence of the $m^{N}$ stick lengths (not distinct) is $O(N^{\delta(-1/(\kappa_{0}-1)+\epsilon^{\prime})})$ for some $\delta>0$ and for every $\epsilon^{\prime}>0$.
In [1], Becker et al. also looked into other variants of the stick fragmentation model, in particular, the Unrestricted 1-Dimension Decomposition Model, where the proportion cuts are independently drawn from the same unspecified probability density $f$. They were able to establish a sufficient condition in terms of the Mellin transforms of the probability density function $f$ for strong Benford behavior to emerge from the stick length distribution. This direction was generalized by Durmić and Miller to a higher dimensional model called the box fragmentation model in [10], and strong Benford behavior was established for the volume of a fragmented box of arbitrary dimension $m$ under a similar condition on the Mellin transforms of $f$. In [6], Betti et al. further generalized the work of Durmić and Miller on box fragmentation model and considered the volume of the lower $d$-dimensional faces of a fragmented box of arbitrary dimension $m$, where $d\leq m$. To state their results, we start with some definitions.
Definition 6.
@@
A set $\mathfrak{B}\subset\mathbb{R}^{m}$ is said to be an $m$-dimensional box if it is a set of the form $a_{1},b_{1}]\times\cdots\times[a_{m},b_{m}]\subset\mathbb{R}^{m}$, where $a_{i}<b_{i}$ are finite numbers.
Definition 7.
@@
$\textup{Vol}_{d}(\mathfrak{B})\ :=\ 2^{m-d}\sum_{|I|=d}\prod_{i\in I}(b_{i}-a_{i}),$ (8)
@@
[6] established a sufficient condition for strong Benford behavior to emerge (see their Theorem 1.9), which involves the volume of the largest $d$-dimensional face.
Theorem 3 (Maximum Criterion).
@@
[6] Let $\mathfrak{B}:=\mathfrak{B}_{0}$ be a fixed $m$-dimensional box and $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\cdots$ be a linear-fragmentation process with proportion cuts $P_{i}$ having probability density functions $f_{i}:(0,1)\rightarrow(0,\infty)$. Let
$V_{d}^{(N)}\ :=\ \textup{Vol}_{d}(\mathfrak{B}_{N})$ (9)
@@
be the sequence of volumes obtained from this process and $\mathfrak{m}_{d}^{(N)}$ denote the maximum product of $d$ sides of $\mathfrak{B}_{N}$. If $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$, then $V_{d}^{(N)}$ also converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$.
In the same paper, the Maximum Criterion was verified for $d=1$ in a special case.
Theorem 4.
@@
Let $P_{i}^{(j)}$ be i.i.d. with $\log_{B}P_{i}^{(j)}$’s having uniform distributions. When $d=1$, the maximum side length of each box
$\mathfrak{m}_{1}^{(N)}\ =\ \textup{max}_{1\leq i\leq m}P_{i}^{(1)}\cdots P_{i}^{(N)}$ (10)
converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$.
@@
It was conjectured in the same paper (see Conjecture 4.1) that for any linear fragmentation process, $m_{d}^{(N)}$ converges to strong Benford behavior.
Conjecture 1.
@@
We prove the conjecture for any $m\geq 2$ and $d\leq m$ under some conditions on the linear fragmentation process, which we shall explain more in Section 3.
Theorem 5.
@@
Suppose that $\mathfrak{B}_{0}\supset\mathfrak{B}_{1}\supset\mathfrak{B}_{2}\supset\cdots$ is a linear fragmentation process on an $m$-dimensional box for an arbitrary $m\geq 2$, such that the proportion cut $P_{i}^{(j)}$ at every stage satisfies the following.
  1. 1.
    Condition 1: $P_{i}^{(j)}$’s are i.i.d. for all $1\leq i\leq m$ and $1\leq j\leq N$. That is, $P_{i}^{(j)}$’s are all independent and there exists a random variable $P$ such that $P_{i}^{(j)}\sim P$ for all $i,j$;
  2. 2.
    Condition 2: Let the mean and the variance of $\log_{B}(P)$ be $\mu_{P}$ and $\sigma_{P}$. Then $\log_{B}(P)$ has finite centered moments, i.e., $\mathbb{E}[(\log_{B}(P)-\mu_{P})^{k}]<\infty$ for all $k\geq 1$ and the characteristic function of $(P-\mu_{P})/\sigma_{P}$ admits Taylor expansion. Moreover, the variance is nonzero;
  3. 3.
    Condition 3: Let $f(x)$ be the probability density function of $\log_{B}(P)$, then we have $\textup{ess sup}_{x}f(x)<\infty$;
  4. 4.
    Condition 4: $f(x)$ is differentiable with $\textup{sup}_{x}|f^{\prime}(x)|<\infty$.
@@
Then this linear fragmentation process satisfies the condition of the Maximum Criterion for any $d\leq m$, i.e., $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$ for any $d\leq m$, thus by Theorem 3, $V_{d}^{(N)}$ converges to strong Benford behavior in base $B$ as $N\rightarrow\infty$ for any $d\leq m$.
2 Proof of Theorem 2
2.1 Preliminary